OpenStudy (anonymous):

what equation would provide the best estimate of 32% of 44.7

OpenStudy (anonymous):

So "of" in math is a hint for multiplication. Therefore, we would use the decimal form of 32%, which is 0.32, and multiply it by 44.7 equation: 0.32*44.7

OpenStudy (texaschic101):

its asking for an estimate, not the exact answer I would guess .30 x 45

OpenStudy (anonymous):

well, if you insist on estimating, 0.33 is approximately 1/3, and is much closer to 0.32. Therefore, you would round 44.7 to 45, and 0.32 to 1/3. Your final equation is 45*1/3 or 45/3. Enjoy.

OpenStudy (anonymous):

\[0.33\approx 1/3, (read 0.3 repeating)\]

OpenStudy (texaschic101):

if this is multiple choice, can you list the answer choices ?

OpenStudy (anonymous):

1/3*40=n 2/3*45=n 1/3*45=n 2/3*44=n

OpenStudy (anonymous):

Thank you very much.

OpenStudy (texaschic101):

1/3 x 45

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Again, the answer I already posted.